OpenStudy (anonymous):

Solve, using the substitution method. 3x + 2y = 11 x – 2 = – 4y

OpenStudy (anonymous):

From 2nd equation, you have x = -4y + 2. Substitute this for x in the first equation. You get: 3(-4y + 2) + 2y = 11 -12y + 6 + 2y = 11 -10y + 6 = 11 -10y = 11-6 -10y = 5 y = -5/10 y = -1/2 Now, find x as -4y + 2 = +2+2 = 4

OpenStudy (anonymous):

Here Are Anwser Choices

OpenStudy (anonymous):

D